# Comparative validation of surgical phase recognition, instrument keypoint estimation, and instrument instance segmentation in endoscopy: Results of the PhaKIR 2024 challenge

### v1

@@ 9 Conclusion
To overcome the limitations of existing datasets and earlier challenges, we organized the PhaKIR challenge – covering surgical surgical phase recognition, keypoint estimation, and instrument instance segmentation – as part of the EndoVis event at MICCAI 2024. For this purpose, we introduced a novel multi-center dataset featuring complete real-world laparoscopic cholecystectomy sequences from 13 patients, annotated consistently for all three tasks. Using this dataset, we systematically benchmarked the performance of submitted algorithms and conducted an in-depth evaluation across a range of metrics. Our findings reveal that surgical phase recognition and instrument instance segmentation currently receive greater attention in the research community than the more complex and underexplored task of instrument keypoint estimation.

### v2

@@ 9 Conclusion
To overcome the limitations of existing datasets and earlier challenges, we organized the PhaKIR challenge – covering surgical phase recognition, keypoint estimation, and instrument instance segmentation – as part of the EndoVis event at MICCAI 2024. For this purpose, we introduced a novel multi-center dataset featuring complete real-world laparoscopic cholecystectomy sequences from 13 patients, annotated consistently for all three tasks. Using this dataset, we systematically benchmarked the performance of submitted algorithms and conducted an in-depth evaluation across a range of metrics. Our findings reveal that surgical phase recognition and instrument instance segmentation currently receive greater attention in the research community than the more complex and underexplored task of instrument keypoint estimation.